OpenStudy (anonymous):

7/9 divided by 21/27

OpenStudy (anonymous):

1

OpenStudy (lgbasallote):

\[\huge \frac{\frac{7}{9}}{\frac{27}{21}} = \frac{7}{9} \times \frac{27}{21}\]

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {a \over b} \div {x \over y} = {a \over b} \times {y \over x}}\)

OpenStudy (anonymous):

You turn the division into multiplication so the second fraction flips.

OpenStudy (anonymous):

Choose the ratio that is equivalent to this one. 1/7 a.3/12 b.14/2 c.4/11 d.3/14

OpenStudy (anonymous):